The mean height of males 20 years or older is 68 inches with a standard deviation of 3.5 inches. The mean height of females 20 years or older is 62 inches with a standard deviation of 2.5 inches. While a male has a height z-score of -1.3525, a female has a height z-score of 1.3525 How tall is he in inches? How tall is she in inches? 25. 26.
Answer #1

25) corresponding height of male =mean+z*std deviation=68-1.3525*3.5=63.2663 inches

26)corresponding height of female =mean+z*std deviation=62+1.3525*2.5=65.3813 inches